y=2x-7
3x-4y=8
help please =]

Answer by elima(1433)

Since you did not give any instructions on how to solve, I will go ahead and solve it by substitution;
y=2x-7
3x-4y=8
Take equation #1 and substitute in equation #2 for y;
3x-4(2x-7)=8
3x-8x+28=8
-5x=8-28
-5x=-20
x=4
Now plug that into equation #1 and solve for y;
y=2(4)-7
y=8-7
y=1
so you have; x=4, y=1
:)